Home‎ > ‎Reports‎ > ‎Crew100A‎ > ‎

day07


Crew 100A Log Book for February 18, 2011

Photos of the day





Commander's Check-In Report

Ivo Ferreira Reporting

Crew Physical Status: Good

Time Departed/Returned from EVA  #6 : 17:00 - 19:00

Brief Narrative of Field Mission Results: We continued with the habitability experiments during the morning. Two of the crew members had a last try on the search and rescue mission of the lost parrot, unfortunately with no success. The day was mostly passed preparing the complex EVA evolving the parrot UAV (the backup one), the Max Rover and a human EVA. A portable relay station was built to extend the capabilities of the WiFi network to be able to control the rover from the Habitat to the mission objective (Coordinates: 38.40617-110.791054, File name: MDRS_Friday_RidgewithJon 183). After a series of problems related to the quality of the signal, directionality of antenna and limited battery status, the objectives of the mission were delayed until late afternoon at which point we were able to locate and recognize (through pictures) the target with the parrot UAV and the Rover. The mission will continue tomorrow with the Human EVA to collect samples at the same location and analyze them for physical and chemical properties back at the Habitat.

EVA Data/Interpretations: It is important to have real time communication between the different actors in an EVA to be able to improve the capabilities of human with UAVs and robots. It is often necessary to work with relay station that have to be tested ahead of time, and positioned in a strategic position (top of mountain) with direct line of sight of both the habitat and the controlled object. Consider using the rover to act as a portable relay station for the UAV signal.

Engineering/Hab Maintenance: Flat tire of the Opportunity ATV, it seems to be leaking quite slowly. Water in the tanks of the GreenHab seems to have evaporated a lot.

Plans for Tomorrow: Finish with EVA to same location of today with astronauts collecting samples for analysis at the laboratory of the Habitat. Continue with experiments: habitability, and EVA to test stress level in the afternoon.

Reports scheduled for today: Engineering, Commander, Journalist

Inventory: None

Support Requested: Refill water tank in trailer and perhaps tire of ATV

Miscellaneous: Yesterday we passed the Hughes network internet limit so we a have limited bandwidth until around 23:00


Commander's Report

Ivo Ferreira Reporting

Everything is getting together now.

After our daily morning routine: temperature pills, Holter monitors and equipments, we spent the morning arranging the last operational details for the forthcoming EVA. Jeffrey had a marathon of tests with Ayako including color distinction and mirror copying, designed to stimulate our brain. Unfortunately, since the satellites cannot live without Jeffrey (or is it the other way around?) he will have to leave the Habitat tomorrow to go back to Germany. We, Max and the parrots are going to miss him a lot!

After lunch Kévin and Romain went on a last search and rescue mission for the missing parrot. We now have to declare it officially lost in battle...

Back in the Habitat, we set up a portable relay station (AKA Christmas tree) with 2 high gain antennas and 2 amplifiers, designed to get our signal from here to Hanksville but forgot about one tiny detail... the battery. Unfortunately the one we had prepared could not give us more that 30 minutes of lifetime powering all this equipment so we had to scavenge the junkyard for a old one with more capacity.

Changed the battery and used our handyman skills to put all the equipment in a box that we placed in a strategic position between the Habitat and the objective and it was time for a test. Unfortunately it was already 5pm and the Sun was almost at the Horizon but we just had to try it... Set up Max and the parrot and there we went... Even though Max returned with the help of a flashlight, it was a complete success! We got some really nice pictures of the site which we are currently analyzing to decide if we will go there tomorrow on an EVA.

Tonight is Jeffrey's last night. We are going to "celebrate" with some non-cooking food (AKA dehydrated beef and vegetable).


Engineering Report

Kévin Theatre Reporting

Generator/Electricity:

· Kitty (Diesel Generator) (Used? Yes/No): Yes

· Honey (Gas Generator) (Used? Yes/No): No

· Battery Charge Level: Checked by DG

o 10.59 V everstart

o 12.37 V armor plate

o 13.32 V green


Notes/Comments (include how many generator runs, approximately when, and any times you needed to turn the generator on/off):

1 generator run, all is good


Fuel Status:

· Diesel (Full/0.75/0.5/0.25/Empty): 0.80

· Propane Tank (Full/0.75/0.5/0.25/Empty): 0.56

· Gasoline Tank (Full/0.75/0.5/0.25/Empty): 11 Gallons

Notes/Comments:


Oil:

· Kitty Oil Quantity (# of Quarts in storage): Leave it to DG

· Honey Oil Quantity (# of Quarts in storage): Leave it to DG

· ATV Oil Quantity (# of Quarts in storage): 6 quarts on the engineering work bench

· Car Oil Quantity (# of Quarts in storage):


Water Status:

· Outside Potable Water Tank Level (inches from bottom): 18

· Trailer Potable Water Tank Level (inches from bottom): 1/10 Full

· Hab Potable Water Tank Level (inches from bottom): 9,5

· Potable Water Meter Reading: 422484


Notes/Comments:


GreenHab:

· Condition of plants in Tank 1: green water

· Condition of plants in Tank 2: green things

· Condition of plants in Tank 3: green things

· Greenhab notes/ comments: The level of water of the 3 tanks has reduced a little


Observatory:

· Telescope Used? (Yes/no) No

Observatory Notes:


Transportation:

Hab Car (NOT RENTAL)

· Used (yes/no): No

· Oil Checked (yes/no): No

· Tire Status: Not checked

· Odometer at end of day: 123611.6

· Notes/Comments on Hab Car:


ATV 1 (Opportunity):

· Used: No

· Oil Checked: no

· Fuel Consumed: Replaced

· Tire Status: 1 tire is flat. We will check this tomorrow.

· Comments:


ATV 2 (Spirit):

· Used: No

· Oil Checked: No

· Fuel Consumed:

· Tire Status:

· Comments: Still not fixed


ATV 3 (Viking I):

· Used: Yes

· Oil Checked: no

· Fuel Consumed: Replaced

· Tire Status: Filled

· Comments:


ATV 4 (Viking II):

· Used: yes

· Oil Checked: no

· Fuel Consumed: Replaced

· Tire Status: Filled

· Comments:


Heating and Ventilation:

· Thermostat setting upstairs (Farenheit): 70

· Actual temperature upstairs (Farenheit): 71

· Thermostat setting downstairs (Farenheit): 65

· Actual temperature downstairs (Farenheit): 64


Computers/Networking Infrastructure:

Notes/comments on internet/computers?: We have consumed yesterday more than we were allowed, so we have a bad connection today.

General Engineering

1. Notes/Comments:


Chef's Report

Ayako Ono and Audrey Lan San Reporting

1a) Today was

- a non-cooking day

1b) Was today a special day celebrated at the main meal with special food or activity?

No.

2) List the foods served at today's main meal (usually the evening meal), giving full names of each.

Alpine Aire Mushroom Pilaf, cheddar broccoli soup, carrot and green beans dehydrated and iced tea from instant mix.

3) List any main meal foods not finished by the crew (leftovers stored for later use or discarded) Example: 2 half servings plate of Mushroom pilaf and 2 plates of carrots and green beans.

There was too much food.

4) Recipes for anything more complicated than following package directions exactly, or rehydrating.

N/A

5) Lessons learned (if any)

Do a better estimation of what is needed for a meal

6) Comments/questions for Kim and Jean, the food study investigators

- any shortages or excesses of food supplies or equipment? (includes "cannot find...", “we are running out of…” “we have too much…” )

- any problems with supplied food? (e.g. spoilage, insect or rodent damage, off flavors, difficulty rehydrating)

- any problems with kitchen equipment, including cleaning supplies?

We miss some hooks to hang the dish towels.

- any cooking questions you need help with?

- any unusual events? (cleaned the oven, spilled all the vanilla, took inventory)

- any other thoughts or suggestions?


Journalist Report

Romain Benchenafi Reporting

We managed to make the drone and the Rover work together for mission TIGER. They have well achieved their objectives and thanks to them the mission was a complete success! That was our seventh day in the MDRS and the last for Jeffrey. We will miss him...